A car of mass 875 kg is traveling 30.0 m/s when the driver applies the brakes, which lock the wheels. The car skids for 5.60 s in the positive -direction before coming to rest. (a) What is the cars acceleration? (b) What magnitude force acted on the car during this time? (c) How far did the car travel? [Ans: ()-536 m/s (b) 4690 N(c) 84,0 m]
Answer #1

(a)

a=\frac {v-u}{t}=\frac {0-30}{5.60}\: m/s^2= -5.36\: m/s^2

(b)

F=ma=(875\times 5.36)\: N = 4690\: N

(c)

s=\frac {v^2-u^2}{2a}=\frac {0^2-30^2}{2(-5.36)}\: m = 84\: m